# Zero-Shot Conditioning of Score-Based Diffusion Models by Neuro-Symbolic Constraints

### v2

@@ 4 Method
where $g(t):[0,1]\rightarrow[0,1]$ satisfies $g(0)=1$ and $g(1)=0$. Sampling from the target distribution then reduces to substituting the score of the base model with the modified score $\tilde{\mathbf{s}}_{c}(\mathbf{x},t)$. Notice that this approach does not require any re-training of the model. The only necessary ingredients are the unconditional score model $\mathbf{s}(\mathbf{x},t)$ and the differentiable constraint $c(\mathbf{x})$ encoding the degree of satisfaction of the desired property.
@@
The difference in our definition with respect to LogLTN is in the logical disjunction ($\lor$): they define it using the LogMeanExp (LME) operator, an approximation of the maximum that is numerically stable and suitable for derivation. They do it at the price of losing the possibility to reduce formulas to the NNF exactly (using De Morgan’s laws) that follows from having as disjunction the dual t-conorm of the conjunction. We choose instead to use the log-probabilistic sum as the disjunction, since in the domain of our experiments it proved numerically stable and effective. In Appendix E we show how to implement the logical disjunction in a numerically stable way.
@@
Score-based models have been mainly used for image generation, adapting them to tabular data and time series requires special care. In particular, the challenge is to correctly model the noise of the target distribution, implying the tricky task of estimating the score at $t\approx 0$ (no noise). We improved the score estimate mainly by parametrizing carefully the score network and using large batches. We provide more details in Appendix B. Correctly estimating the score at $t\approx 0$ is fundamental to make our method work in practice, since it allows us to perform Langevin MCMC at $t\approx 0$, where the conditional score is known without approximation. Combining a correct score estimation at $t\approx 0$ with many steps of Langevin MCMC allows us to (asymptotically) sample from the exact conditional distribution, in particular when the data distribution is not particularly multimodal.
@@ 5 Experiments
We mostly used unconditional models based on denoising score matching and SDEs, following closely Song et al. (2021). Conditional generation hyperparameters and more details about models are described in Appendix D and A.
@@
where $\text{class}(\mathbf{x},i):=P\{\mathbf{x}\text{ is classified as }i\}=1$, and $P$ is obtained from a pre-trained classifier. The generated pairs of digits add up to ten in $\approx 96\%$ of cases444according to classes assigned by the classifier, however, only 2-8 and 4-6 pairs were generated. More details are discussed in Appendix G.
@@
In general, we observed the effectiveness of our method to depend on the task. While for some experiments keeping a large constraint strength $k$ (as $k=30$) was sufficient, for other experiments we needed to tune $k$ until a satisfactory trade-off between constraint satisfaction and quality of samples was reached. In Appendix G we discuss more experiments and cases where this trade-off was poorer.
@@
On the other side, we noticed a sensible improvement in some of our image-based conditional generation tasks. For example, relatively to the MNIST digits’ sum experiment, universal guidance allowed us to generate all possible pairs, that always summed up to ten. We infer that universal guidance can be useful when the objective is to generate high-quality samples that satisfy a given constraint, but it can lead to large biases when modelling the conditional distribution, that is of primary interest in some contexts, as with tabular data and time series. Moreover, with these experiments we verified that our Log-probabilistic logic is effective also within other guidance techniques. We discuss the experiments more deeply in Appendix F.
@@ 6 Conclusion
Appendix A Unconditional models
@@
Appendix B Training a score model for tabular data
@@ Appendix C Normalization
Appendix D Constrained generation settings
@@
Appendix E Numerically stable implementation of logical disjunction
@@
Appendix F Universal guidance comparison experiments
@@
Appendix G Other experiments
@@
Appendix H Hardware specifications

### v3

@@ 3 Related Work
Despite the existence of many related prior works with different focuses, our goal is different and more challenging: obtaining samples distributed according to the target conditional distribution, while previous methods rather aim at obtaining high quality samples. To our knowledge, there are no works focusing on the correct approximation of the conditional distribution for tabular data as we do. The best comparison we can do is with Bansal et al. (2023), since it is the state of the art for diffusion based zero-shot conditional generation and a synthesis of previous techniques. We show that our method is significantly better with tabular data when the objective is the approximation of the conditional distribution.
@@ 4 Method
where $g(t):[0,1]\rightarrow[0,1]$ satisfies $g(0)=1$ and $g(1)=0$. This is equivalent to extending the domain of the constraint to noisy data points $c(\mathbf{x},t)$ and then approximating it with $c(\mathbf{x},t)=g(t)c(\mathbf{x})$. Sampling from the target distribution then reduces to substituting the score of the base model with the modified score $\tilde{\mathbf{s}}_{c}(\mathbf{x},t)$. Notice that this approach does not require any re-training of the model. The only necessary ingredients are the unconditional score model $\mathbf{s}(\mathbf{x},t)$ and the differentiable constraint $c(\mathbf{x})$ encoding the degree of satisfaction of the desired property.
@@
The difference in our definition with respect to LogLTN is in the logical disjunction ($\lor$): they define it using the LogMeanExp (LME) operator, an approximation of the maximum that is numerically stable and suitable for derivation. They do it at the price of losing the possibility to reduce formulas to the NNF exactly (using De Morgan’s laws) that follows from having as disjunction the dual t-conorm of the conjunction. We choose instead to use the log-probabilistic sum as the disjunction, since in the domain of our experiments it proved numerically stable and effective. In Appendix F we show how to implement the logical disjunction in a numerically stable way.
@@
Score-based models have been mainly used for image generation, adapting them to tabular data and time series requires special care. In particular, the challenge is to correctly model the noise of the target distribution, implying the tricky task of estimating the score at $t\approx 0$ (no noise). We improved the score estimate mainly by parametrizing carefully the score network and using large batches. We provide more details in Appendix C. Correctly estimating the score at $t\approx 0$ is fundamental to make our method work in practice, since it allows us to perform Langevin MCMC at $t\approx 0$, where the conditional score is known without approximation. Combining a correct score estimation at $t\approx 0$ with many steps of Langevin MCMC allows us to (asymptotically) sample from the exact conditional distribution, in particular when the data distribution is not particularly multimodal.
@@ 5 Experiments
We mostly used unconditional models based on denoising score matching and SDEs, following closely Song et al. (2021). Conditional generation hyperparameters and more details about models are described in Appendix E and B.
@@
where $\text{class}(\mathbf{x},i):=P\{\mathbf{x}\text{ is classified as }i\}=1$, and $P$ is obtained from a pre-trained classifier. The generated pairs of digits add up to ten in $\approx 96\%$ of cases444according to classes assigned by the classifier, however, only 2-8 and 4-6 pairs were generated. More details are discussed in Appendix H.
@@
In general, we observed the effectiveness of our method to depend on the task. For most experiments involving tabular data, keeping a large constraint strength $k$ (as $k=30$) was sufficient. However, for some experiments, as most tasks involving images, we needed to tune the constraint strength until a satisfactory trade-off between constraint satisfaction and quality of samples was reached. In Appendix H we discuss more experiments and cases where this trade-off was poorer.
@@
On the other side, we noticed a sensible improvement in some of our image-based conditional generation tasks. For example, relatively to the MNIST digits’ sum experiment, universal guidance allowed us to generate all possible pairs, that always summed up to ten. We infer that universal guidance can be useful when the objective is to generate high-quality samples that satisfy a given constraint, but it can lead to large biases when modelling the conditional distribution, that is of primary interest in some contexts, as with tabular data and time series. Moreover, with these experiments we verified that our Log-probabilistic logic is effective also within other guidance techniques. We discuss the experiments more deeply in Appendix G.
@@ 6 Conclusion
Acknowledgements
This research was supported by Aindo, which has funded the PhD of the first author and provided the computational resources. Additionally, this study was carried out within the PNRR research activities of the consortium iNEST funded by the European Union Next-GenerationEU (PNRR, Missione 4 Componente 2, Investimento 1.5 – D.D. 1058 23/06/2022, ECS_00000043).
@@ Acknowledgements
Appendix A Comparison with Logic Tensor Networks
Deep learning allows to learn complex statistical properties at the cost of data inefficiency, struggle at generalization, lack of comprehensibility and ignorance of prior knowledge. In contrast, Symbolic AI has proven useful at several reasoning tasks (as theorem proving, logical inference, verification) but at the cost of high computational complexity and struggle with large and inaccurate data. Neurosymbolic AI aims at combining the strength of both paradigms. Recent research has focused on developing differentiable approaches for knowledge representation and reasoning. This includes relaxing logical operators into continuous operations using fuzzy semantics. Logic Tensor Networks (Badreddine et al. 2020) is an example of such a neurosymbolyc framework, that supports learning and reasoning about data with a knowledge base. In LTN, one can represent and perform several tasks of deep learning (as clustering, classification, embedding learning) with a fully differentiable first order logic language, that they call Real Logic, assuming truth values in the interval [0,1]. Real Logic employs fuzzy semantics, where logical connectives ($\land$, $\lor$, …) and quantifiers ($\forall$, $\exists$) are mapped to fuzzy operators like t-norms and aggregators. These operators enable formulas to have degrees of truth in the interval [0, 1], which are optimized during training by maximizing the satisfaction of a knowledge base of logical rules. In LTN the preferred fuzzy logic is the Stable Product Real Logic. LogLTN (Badreddine, Serafini, and Spranger 2023) is a recent extension aiming to improve numerical stability in fuzzy semantics by working in logarithm space, as maximizing the log satisfaction of a formula is equivalent to maximizing its satisfaction.
Our LogProbabilistic logic has many similarities with the fuzzy logic used in LogLTN. We devolped it independently mainly motivated by probabilistic principles, as our formulas can be seen as describing un-normalized distribution encoding prior knowledge. In practice, the difference with their approach lies in the mapping of the logical disjunction. While we used the probabilistic disjunction as shown in table 1, they used a surrogate of the maximum function, the LogMeanExp operator, as the maximum function is single-passing (the gradient on passed through one of the arguments):
$\operatorname{LME}(\mathbf{x}\mid\alpha,C)=\frac{1}{\alpha}\left(C+\log\left(\frac{\sum_{i=1}^{n}e^{\alpha x_{i}-C}}{n}\right)\right)$
where $C=\max_{i}(\alpha x_{i})$ and $\alpha$ is an hyperparameter. Authors claim that the LME operator is numerically stable, well-bounded, and suitable for derivation. Our numerically stable implementation of the logical disjunction that we discuss in Appendix F performed well in our experiments, being at the same time well grounded in probabilistic principles.
Appendix B Unconditional models
@@
Appendix C Training a score model for tabular data
@@ Appendix D Normalization
Appendix E Constrained generation settings
@@
Appendix F Numerically stable implementation of logical disjunction
@@
Appendix G Universal guidance comparison experiments
@@
Appendix H Other experiments
@@
Appendix I Hardware specifications